Calculate the charged stored on a 3-pF capacitor with 20V across it


1
Expert's answer
2020-08-03T15:10:59-0400

The charge stored on a capacitor

"q=CV\\
=(3\times 10^{-12}\:\rm F)\times (20\:\rm V)=60\times 10^{-12}\:\rm C=60\: pC"